OpenStudy (anonymous):

Will become Fann and give medal to best answer:) thanx <3

OpenStudy (anonymous):

Clarissa is having lunch at a sandwhich shop. She can chose white bread or pumpernickel bread. Her options for filling are turkey, tuna, ham, or egg salad. Her choices for condiments are mayonnaise, salad dressing, or mustard. How many different sandwhich choices does clarrisa have? Answer choices: A: 36 B:6 C:24 D:12

OpenStudy (anonymous):

A I think

OpenStudy (anonymous):

Have you tried listing the cases? That is the easiest way to figure out simpler probability questions. In this case there aren't that many.

OpenStudy (anonymous):

alternatively, you can write the options in nCr notation, and multiply. i.e. \[\left(\begin{matrix}2 \\ 1\end{matrix}\right)*\left(\begin{matrix}4 \\ 1\end{matrix}\right)*\left(\begin{matrix}3 \\ 1\end{matrix}\right)\] Either way the results are the same. And not A.